OpenStudy (anonymous):

water freezes at 32"f or 0"c use temperature conversion formulas to show that these to measures are indeed the same temperature.

OpenStudy (anonymous):

\[C * \frac{9}{5} +32 = F\] so if our C=0 then : \[0 * \frac{9}{5} + 32= 0+32=32=F\]

OpenStudy (anonymous):

thankyou :) @hobbs978

OpenStudy (anonymous):

no problem :)